# Aerial Active STAR-RIS-assisted Satellite-Terrestrial Covert Communications

### v2

@@ Abstract
An integration of satellites and terrestrial networks is crucial for enhancing performance of next generation communication systems. However, the networks are hindered by the long-distance path loss and security risks in dense urban environments. In this work, we propose a satellite-terrestrial covert communication system assisted by the aerial active simultaneous transmitting and reflecting reconfigurable intelligent surface (AASTAR-RIS) to improve the channel capacity while ensuring the transmission covertness. Specifically, we first derive the minimal detection error probability (DEP) under the worst condition that the Warden has perfect channel state information (CSI). Then, we formulate an AASTAR-RIS-assisted satellite-terrestrial covert communication optimization problem (ASCCOP) to maximize the sum of the fair channel capacity for all ground users while meeting the strict covert constraint, by jointly optimizing the trajectory and active beamforming of the AASTAR-RIS. Due to the challenges posed by the complex and high-dimensional state-action spaces as well as the need for efficient exploration in dynamic environments, we propose a generative deterministic policy gradient (GDPG) algorithm, which is a generative deep reinforcement learning (DRL) method to solve the ASCCOP. Concretely, the generative diffusion model (GDM) is utilized as the policy representation of the algorithm to enhance the exploration process by generating diverse and high-quality samples through a series of denoising steps. Moreover, we incorporate an action gradient mechanism to accomplish the policy improvement of the algorithm, which refines the better state-action pairs through the gradient ascent. Simulation results demonstrate that the proposed approach significantly outperforms important benchmarks, and also validate the performance and practicality under different algorithm parameters and environment settings.
@@ I Introduction
In such cases, low-altitude economy (LAE) offers new opportunities for the development of satellite-terrestrial communications [Jiahui2025]. As the core supporters of LAE, low-altitude platforms (LAPs), e.g, unmanned aerial vehicles (UAVs), can be deployed as the aerial relay nodes between satellites and ground users. Compared to the fixed relay communication mode, LAP-based aerial relay communications can bring an additional degree of freedom by dynamically adjusting their trajectories, thereby improving the channel capacity for the users in various locations. While the high line-of-sight (LoS) probability of air-to-ground channels improves the transmission performance, the inherent broadcast nature of wireless communications simultaneously heightens the security vulnerabilities [Ding2024]. As such, two widely used conventional techniques for ensuring secure communications are the encryption method in the high-layer network protocol stack and physical layer security [li2024two]. However, these methods cannot fundamentally address security threats of satellite-terrestrial communications, particularly in the scenarios involving advanced technologies such as quantum computers, which could potentially break the conventional encryption methods, and more intelligent adversaries that may adapt to exploit emerging decoding techniques [Chen2023].
@@
I-A Prior Works
We review related works from three aspects, including the architecture of satellite-terrestrial covert communications, uncertainty sources and performance metrics in satellite-terrestrial covert communications, and optimization methods for satellite-terrestrial covert communications, which motivates us to contribute this work.
I-A1 Architecture of Satellite-Terrestrial Covert Communications
Existing works in the architecture of satellite-terrestrial covert communications can be mainly divided into direct mode and relay mode. For the direct mode, Zhang et al. [Zhanglei2024] investigated an ultra-dense low Earth orbit (LEO) satellite downlink system, where the sidelobe leakages of other satellites are utilized to ensure the covertness. In [Jia2025], the authors developed a direct satellite-terrestrial downlink system by employing the rate-splitting multiple access technique to transmit jamming signals. However, this mode typically overlooks the long-distance path loss and complex propagation environments between satellites and ground users, which are particularly problematic in dense urban areas. In terms of the relay mode, Wu et al. [Wu2022] studied the ground device-based satellite-terrestrial covert relay communications, where multiple ground base stations are randomly selected to relay the wireless signal. Moreover, the authors in [Song2023] presented a RIS-based satellite-terrestrial covert relay communication system, where a passive RIS is deployed on the surface of the building to optimize the signal reflection. However, the fixed position of relay platforms limits their adaptability in dynamic environments and reduces the flexibility of the system.
I-A2 Uncertainty Source and Performance Metrics in Satellite-Terrestrial Covert Communications
@@
I-A3 Optimization Methods for Satellite-Terrestrial Covert Communications
Existing studies on optimization methods for satellite-terrestrial covert communication systems primarily relied on convex optimization theory. For example, Hui et al. [Hui2024] proposed a two-stage iterative method with the block coordinate descent to maximize the average detection error probability (DEP). In [Cao2024], the authors utilized convex optimization to minimize the average age-of-information by taking the derivative of the transmission power of the transmitter in satellite-terrestrial covert communications. Moreover, Jia et al. [Jia2025] maximized the minimum covert rate by using the fractional programming, semidefinite relaxation and the S-procedure. However, these methods often assume static environments or require the channel state information (CSI) of the Warden. These assumptions limit the practicality of these methods in satellite-terrestrial covert communication systems, where the instantaneous CSI of the Warden is difficult to obtain and the detection threshold of the Warden may evolve over time.
TABLE I: Comparison Between Our Work and Existing Works
Reference
(2022-2025)
Satellite-Terrestrial
Covert Communications
Uncertainty
Sources
Performance Metrics Decision Variables Optimization Methods
Relay
Mode
RIS-assisted
Relay Mode
Movable RIS-assisted
Relay Mode
Environmental
Noise
Channel
Capacity
Fairness Trajectory
Active
Beamforming
DRL
Generative
Models
@@ I-B Motivations and Contributions
  • AASTAR-RIS-assisted Satellite-Terrestrial Covert Communication Systems: We propose a novel AASTAR-RIS-assisted satellite-terrestrial covert communication system that leverages an LAP equipped with active STAR-RIS to act as a relay between a geosynchronous Earth orbit (GEO) satellite and the ground users. This is particularly beneficial in dense urban scenarios where the direct link is limited. Specifically, the system utilizes the flexibility of the LAP and the reconfigurable capability of an active STAR-RIS to dynamically adjust wireless channel properties, thereby enhancing the transmission rate by the signal amplification and concentration while ensuring the covert transmission by manipulating the electromagnetic wave propagation.
@@
I-C Organization and Notations
The remainder of this paper is organized as follows. We first describe the system model in Section II. In Section III, the covert requirement of the system is derived and ASCCOP is formulated and analyzed. Next, we present a generative DRL algorithm to solve the proposed ASCCOP in Section IV. Section V shows numerical results and provides corresponding discussions. Finally, Section VI concludes this paper.
@@ II-D Communication Model
II-E Binary Hypothesis Testing at Warden
@@
In the binary testing problem, the decision rule of the Warden can be expressed as follows [Wang2025]:
@@
III Problem Formulation and Analysis
@@ IV-A3 Reward
IV-B Conventional Deterministic Policy Gradient Algorithm
@@ IV-B2 Policy Improvement
IV-C Generative Deterministic Policy Gradient Algorithm
IV-C1 Motivation of Adopting GDM
@@ 11         Sample a batch from $\mathcal{D}$;
IV-C3 Action Gradient for Policy Improvement
@@
IV-D Overall Structure and Complexity Analysis
@@ V Performance Evaluation
V-A Simulation Setup
V-A1 Environment, Platform and Model Details
@@ V-A2 Benchmarks
V-B Simulation Results
V-B1 Comparison with Other Approaches
@@
V-B2 Comparison with State-of-the-Art DRL Algorithms
@@
V-B3 Impact of Different Diffusion Steps
@@
V-B4 Impact of Different Learning Rates
@@
V-B5 Impact of Different Covert Requirements
@@
V-B6 Impact of Different AASTAR-RIS Sizes

### v3

@@ Abstract
An integration of satellites and terrestrial networks is crucial for enhancing performance of next-generation communication systems. However, the networks are hindered by the long-distance path loss and security risks in dense urban environments. In this work, we propose a satellite-terrestrial covert communication system assisted by the aerial active simultaneous transmitting and reflecting reconfigurable intelligent surface (AASTAR-RIS) to improve the channel capacity while ensuring the transmission covertness. Specifically, we first derive the minimal detection error probability (DEP) under the worst condition that the Warden has perfect channel state information (CSI). Then, we formulate an AASTAR-RIS-assisted satellite-terrestrial covert communication optimization problem (ASCCOP) to maximize the sum of the fair channel capacity for all ground users while meeting the strict covert constraint, by jointly optimizing the trajectory and active beamforming of the AASTAR-RIS. Due to the challenges posed by the complex and high-dimensional state-action spaces as well as the need for efficient exploration in dynamic environments, we propose a generative deterministic policy gradient (GDPG) algorithm, which is a generative deep reinforcement learning (DRL) method to solve the ASCCOP. Concretely, the generative diffusion model (GDM) is utilized as the policy representation of the algorithm to enhance the exploration process by generating diverse and high-quality samples through a series of denoising steps. Moreover, we incorporate an action gradient mechanism to accomplish the policy improvement of the algorithm, which refines the better state-action pairs through the gradient ascent. Simulation results demonstrate that the proposed approach significantly outperforms important benchmarks, and also validate the performance and practicality under different algorithm parameters and environment settings.
@@ 1 Introduction
In such cases, low-altitude economy (LAE) offers new opportunities for the development of satellite-terrestrial communications [Jiahui2025]. As the core supporters of LAE, low-altitude platforms (LAPs), e.g, unmanned aerial vehicles (UAVs), can be deployed as the aerial relay nodes between satellites and ground users. Compared to the fixed relay communication mode, LAP-based aerial relay communications can bring an additional degree of freedom by dynamically adjusting their trajectories, thereby improving the channel capacity for the users in various locations [Hu2021]. While the high line-of-sight (LoS) probability of air-to-ground channels improves the transmission performance, the inherent broadcast nature of wireless communications simultaneously heightens the security vulnerabilities [Ding2024]. As such, two widely used conventional techniques for ensuring secure communications are the encryption method in the high-layer network protocol stack and physical layer security [li2024two]. However, these methods cannot fundamentally address security threats of satellite-terrestrial communications, particularly in the scenarios involving advanced technologies such as quantum computers, which could potentially break the conventional encryption methods, and more intelligent adversaries that may adapt to exploit emerging decoding techniques [Chen2023].
@@
1.1 Prior Works
We review related works from three aspects, including the architecture of satellite-terrestrial covert communications, uncertainty sources and performance metrics in satellite-terrestrial covert communications, and optimization methods for covert communications, which motivates us to contribute this work.
1.1.1 Architecture of Satellite-Terrestrial Covert Communications
Existing works in the architecture of satellite-terrestrial covert communications can be mainly divided into direct mode and relay mode. For the direct mode, Zhang et al. [Zhanglei2024] investigated an ultra-dense low Earth orbit (LEO) satellite downlink system, where the sidelobe leakages of other satellites are utilized to ensure the covertness. In [Jia2025], the authors developed a direct satellite-terrestrial downlink system by employing the rate-splitting multiple access technique to transmit jamming signals. However, this mode typically overlooks the long-distance path loss and complex propagation environments between satellites and ground users, which are particularly problematic in dense urban areas. In terms of the relay mode, Wu et al. [Wu2022] studied the ground device-based satellite-terrestrial covert relay communications, where multiple ground base stations are randomly selected to relay the wireless signal. Moreover, the authors in [Song2023] presented a RIS-based geosynchronous Earth orbit (GEO) satellite-terrestrial covert relay communication system, where a passive RIS is deployed on the surface of the building to optimize the signal reflection. However, the fixed position of relay platforms limits their adaptability in dynamic environments and reduces the flexibility of the system.
1.1.2 Uncertainty Source and Performance Metrics in Satellite-Terrestrial Covert Communications
@@
1.1.3 Optimization Methods for Covert Communications
Existing studies on optimization methods for covert communications primarily relied on convex optimization theory. For example, Hui et al. [Hui2024] proposed a two-stage iterative method with the block coordinate descent to maximize the average detection error probability (DEP). In [Cao2024], the authors utilized convex optimization to minimize the average age-of-information by taking the derivative of the transmission power of the transmitter in satellite-terrestrial covert communications. Moreover, Jia et al. [Jia2025] maximized the minimum covert rate by using the fractional programming, semidefinite relaxation and the S-procedure. Furthermore, Xing et al. [Xing2024] proposed a covert transmission scheme for cell-free Internet-of-Thing networks that leverages the Lagrangian dual algorithm and semidefinite relaxation to jointly optimize beamforming and artificial noise vectors, thereby improving the covert transmit rate. However, these methods often assume static environments or require the channel state information (CSI) of the Warden. These assumptions limit the practicality of these methods in satellite-terrestrial covert communication systems, where the instantaneous CSI of the Warden is difficult to obtain and the detection threshold of the Warden may evolve over time.
TABLE I: Comparison Between Our Work and Existing Works in Satellite-Terrestrial Covert Communications
The relay mode in satellite-terrestrial communication systems is particularly suitable for scenarios where direct satellite-user links are obstructed or suffer from severe path loss, such as in dense urban environments and complex terrains. It enhances the coverage and channel capacity of satellite-terrestrial communication system by introducing an intermediate node (e.g., RIS or aerial platform) to forward signals.
Reference (2022-2025) Satellite-Terrestrial Covert Communications Uncertainty Sources Performance Metrics Decision Variables Optimization Methods
Relay Mode RIS-assisted Relay Mode Movable RIS-assisted Relay Mode Environmental Noise Channel Capacity Fairness Trajectory Active Beamforming DRL Generative Models
@@ 1.2 Motivations and Contributions
  • AASTAR-RIS-assisted Satellite-Terrestrial Covert Communication Systems: We propose a novel AASTAR-RIS-assisted satellite-terrestrial covert communication system that leverages an LAP equipped with active STAR-RIS to act as a relay between a GEO satellite and the ground users. This is particularly beneficial in dense urban scenarios where the direct link is limited. Specifically, the system utilizes the flexibility of the LAP and the reconfigurable capability of an active STAR-RIS to dynamically adjust wireless channel properties, thereby enhancing the transmission rate by the signal amplification and concentration while ensuring the covert transmission by manipulating the electromagnetic wave propagation.
@@
1.3 Organization and Notations
The remainder of this paper is organized as follows. We first describe the system model in Section 2. In Section 3, the covert requirement of the system is derived and ASCCOP is formulated and analyzed. Next, we present a generative DRL algorithm to solve the proposed ASCCOP in Section 4. Section 5 shows numerical results and provides corresponding discussions. Finally, Section 6 concludes this paper.
@@ 2.4 Communication Model
2.5 Binary Hypothesis Testing at Warden
@@
In the binary testing problem, the decision rule of the Warden can be expressed as follows [Hieu2023]:
@@
3 Problem Formulation and Analysis
@@ 4.1.3 Reward
4.2 Conventional Deterministic Policy Gradient Algorithm
@@ 4.2.2 Policy Improvement
4.3 Generative Deterministic Policy Gradient Algorithm
4.3.1 Motivation of Adopting GDM
@@ 11         Sample a batch from $\mathcal{D}$;
4.3.3 Action Gradient for Policy Improvement
@@
4.4 Overall Structure and Complexity Analysis
@@ 5 Performance Evaluation
5.1 Simulation Setup
5.1.1 Environment, Platform and Model Details
@@ 5.1.2 Benchmarks
5.2 Simulation Results
5.2.1 Comparison with Other Approaches
@@
5.2.2 Comparison with State-of-the-Art DRL Algorithms
@@
5.2.3 Impact of Different Diffusion Steps
@@
5.2.4 Impact of Different Learning Rates
@@
5.2.5 Impact of Different Covert Requirements
@@
5.2.6 Impact of Different AASTAR-RIS Sizes
@@ 6 Conclusion
[Uncaptioned image] Chuang Zhang received the B.S. degree in computer science and technology from Jilin University, Changchun, China, in 2021, where he is currently pursuing the Ph.D. degree with the College of Computer Science and Technology. He is also a visiting PhD degree student with the Singapore University of Technology and Design, Singapore. His current research interests include low-altitude wireless networks, secure communications, distributed beamforming, and reinforcement learning.
[Uncaptioned image] Geng Sun (Senior Member, IEEE) received the B.S. degree in communication engineering from Dalian Polytechnic University, and the Ph.D. degree in computer science and technology from Jilin University, in 2011 and 2018, respectively. He was a Visiting Researcher with the School of Electrical and Computer Engineering, Georgia Institute of Technology, USA. He is a Professor in the College of Computer Science and Technology at Jilin University. Currently, he is working as a visiting scholar at the College of Computing and Data Science, Nanyang Technological University, Singapore. He has published over 100 high-quality papers, including IEEE TMC, IEEE JSAC, IEEE/ACM ToN, IEEE TWC, IEEE TCOM, IEEE TAP, IEEE IoT-J, IEEE TIM, IEEE INFOCOM, IEEE GLOBECOM, and IEEE ICC. He serves as the Associate Editors of IEEE Communications Surveys & Tutorials, IEEE Transactions on Communications, IEEE Transactions on Vehicular Technology, IEEE Transactions on Network Science and Engineering, IEEE Transactions on Network and Service Management and IEEE Networking Letters. He serves as the Lead Guest Editor of Special Issues for IEEE Transactions on Network Science and Engineering, IEEE Internet of Things Journal, IEEE Networking Letters. He also serves as the Guest Editor of Special Issues for IEEE Transactions on Services Computing, IEEE Communications Magazine, and IEEE Open Journal of the Communications Society. His research interests include Low-altitude Wireless Networks, UAV communications and Networking, Mobile Edge Computing (MEC), Intelligent Reflecting Surface (IRS), Generative AI and Agentic AI, and deep reinforcement learning.
[Uncaptioned image] Jiahui Li received his B.S. in Software Engineering, and M.S. and Ph.D. in Computer Science and Technology from Jilin University, Changchun, China, in 2018, 2021, and 2024, respectively. He was a visiting Ph.D. student at the Singapore University of Technology and Design (SUTD). He currently serves as an assistant researcher in the College of Computer Science and Technology at Jilin University. His current research focuses on integrated air-ground networks, UAV networks, wireless energy transfer, and optimization.
[Uncaptioned image] Jiacheng Wang received the Ph.D. degree from the School of Communication and Information Engineering, Chongqing University of Posts and Telecommunications, Chongqing, China. He is currently a Research Associate in computer science and engineering with Nanyang Technological University, Singapore. His research interests include wireless sensing, semantic communications, and metaverse.
[Uncaptioned image] Ruichen Zhang is currently a Research Fellow with the College of Computing and Data Science, Nanyang Technological University (NTU), Singapore. His research interests include the Internet of Agents, LLM-empowered networking, reinforcement learning-enabled wireless communication, generative AI models, and agentic AI. He has received two Best Paper Awards from IEEE IWCMC and has three ESI Highly Cited Papers and two ESI hot papers. He also serves as a Guest Editor for several IEEE journals, including IEEE TCCN, IEEE TAI, and IEEE JSAS. In addition, he serves as an editor of IEEE TCOM and the Managing Editor of the IEEE TNSE.
[Uncaptioned image] Dusit Niyato (Fellow, IEEE) received the B.Eng. degree from the King Mongkuts Institute of Technology Ladkrabang (KMITL), Thailand, in 1999, and the Ph.D. degree in electrical and computer engineering from the University of Manitoba, Canada, in 2008. He is currently a Professor with the College of Computing and Data Science, Nanyang Technological University, Singapore. His research interests include the Internet of Things (IoT), machine learning, and incentive mechanism design.
[Uncaptioned image] Shiwen Mao (Fellow, IEEE) is a Professor and the Earle C. Williams Eminent Scholar Chair, and the Director of the Wireless Engineering Research and Education Center, Auburn University, Auburn, AL, USA. His research interest includes wireless networks, multimedia communications, and smart grid. He received the IEEE ComSoc MMTC Outstanding Researcher Award in 2023, the IEEE ComSoc TC-CSR Distinguished Technical Achievement Award in 2019, and the NSF CAREER Award in 2010. He is a co-recipient of the 2022 Best Journal Paper Award of IEEE ComSoc eHealth Technical Committee, the 2021 Best Paper Award of Elsevier/Digital Communications and Networks (KeAi), the 2021 IEEE Internet of Things Journal Best Paper Award, the 2021 IEEE Communications Society Outstanding Paper Award, the IEEE Vehicular Technology Society 2020 Jack Neubauer Memorial Award, the 2018 ComSoc MMTCBestJournal Paper Award and the 2017 Best Conference Paper Award, the 2004 IEEE Communications Society Leonard G. Abraham Prize in the Field of Communications Systems, and several ComSoc technical committee and conference best paper/demo awards. He is the Editor-in-Chief of IEEE TRANSACTIONS ON COGNITIVE COMMUNICATIONS AND NETWORKING. He is a Distinguished Lecturer of IEEE Communications Society and the IEEE Council of RFID.
[Uncaptioned image] Abbas Jamalipour (Fellow, IEEE) received the Ph.D. degree in electrical engineering from Nagoya University, Nagoya, Japan, in 1996. He is currently a Professor of ubiquitous mobile networking with The University of Sydney. He has authored nine technical books, eleven book chapters, more than 550 technical papers, and five patents, all in wireless communications and networking. He was a member of the Board of Governors of the IEEE Communications Society. He has been an Elected Member of the Board of Governors of the IEEE Vehicular Technology Society since 2014. He is a member of the Advisory Board of IEEE Internet of Things Journal. He is a fellow of the Institute of Electrical, Information, and Communication Engineers (IEICE), and the Institution of Engineers Australia, an ACM Professional Member, and an IEEE Distinguished Speaker. He was a recipient of a number of prestigious awards, such as the 2019 IEEE ComSoc Distinguished Technical Achievement Award in Green Communications, the 2016 IEEE ComSoc Distinguished Technical Achievement Award in Communications Switching and Routing, the 2010 IEEE ComSoc Harold Sobol Award, the 2006 IEEE ComSoc Best Tutorial Paper Award, and more than 15 best paper awards. He has been the General Chair or the Technical Program Chair of several prestigious conferences, including IEEE ICC, GLOBECOM, WCNC, and PIMRC. He is the Editor-in-Chief of IEEE Transactions on Vehicular Technology. He was the President of the IEEE Vehicular Technology Society from 2020 to 2021. Previously, he held the positions of the Executive Vice-President and the Editor-in-Chief of VTS Mobile World. He was the Editor-in-Chief of IEEE Wireless Communications and the Vice President-Conferences. He sits on the Editorial Board of IEEE Access and several other journals.